# PRDP: Progressively Refined Differentiable Physics

### v1

@@ 2 Differentiating Iterative Linear Solvers
$(\nabla_{\theta}L)^{T}=\bar{\theta}^{T}=\left(\vec{1}^{T}\;\underbrace{{\bm{J}}_{l}|_{\mathcal{P}_{K}(\theta)}}_{\text{primal inacc.}}\right)\;\underbrace{{\bm{J}}_{\mathcal{P}_{K}}|_{\theta}}_{\text{adjoint inacc.}}.$ (2)
@@ 2.2 Scheduling inner iterations
Since the outer problem of equation 1 is solved iteratively, the gradient $\nabla_{\theta}L$ can be coarse at the beginning of the outer iterations and only requires the highest fidelity towards the end (Pedregosa, 2016). This fundamental idea was developed in the domain of hyperparameter optimization, and we demonstrate its application to physical inverse problems. Consider the Poisson equation, which is a prototypical elliptic partial differential equation found in many areas of science and engineering. Most discretization techniques lead to a linear system of equations with $N$ degrees of freedom (for an example discretization, refer to section D.1).
@@ 3 Progressively Refined Differentiable Physics
Figure 5 illustrates the basic concept of the PRDP control algorithm. The exact algorithm is detailed in pseudocode in Algorithm 4. Its implementation in a training pipeline that uses differentiable physics is represented by the should_refine function in listing 3.
{listing}{minted}
[frame=lines, framesep=2mm, fontsize=, ]python def update(model, state_in, state_out, inner_iterations): # neural network gives first time step, physics gives second time step state_1 = model(state_in) state_2 = physics(state_1, inner_iterations) loss, grad = value_and_grad(loss_fn)(state_2, state_out) model = update_model(model, grad) return model, loss
inner_iterations = 1 val_loss_history = [evaluate(model, val_dataloader))] for epoch in range(epochs): for state_in, state_out in dataloader: model, loss = update(model, state_in, state_out, inner_iterations) val_loss_history.append(evaluate(model, val_dataloader)) if should_refine(val_loss_history): inner_iterations += 1 A typical mixed-chain learning pipeline (figure 1) as used in our neural emulator learning experiments. Training begins with coarse physics, using, e.g., 1 solver iteration. The should_refine function applies PRDP, determining when to refine and progressively increasing the inner solver iterations during training.
@@ Appendix A Notation
  • .$^{T}$: Transpose of a matrix or vector.
@@ C.2 Unrolled Differentiaion
[…]
@@ D.1 Poisson Equation in 1D
For a finite difference discretization, the domain $\Omega=(0,1)$ is divided equidistantly into $N+2$ grid points, out of which two are related to the prescribed value on the boundary and can, hence, be ignored. As such, there are $N$ degrees of freedom that make up the discrete solution vector ${\bm{u}}_{h}\in\mathbb{R}^{N}$. Equally, the right-hand side is discretized at the same points and negated, yielding ${\bm{b}}_{h}$. At index $i$ interior to the domain, the second derivative is approximated with the three-point stencil
@@
in which the first matrix is due to the discretized Laplace operator in one dimension. We denote it ~$\tilde{{\mathcal{L}}}_{1}$ (the tilde is to distinguish it from the Laplacian matrices on periodic boundaries of the following sections). The subscript is to indicate the one-dimensional setting. It is tridiagonal and solely defines the system to be solved with ${\bm{A}}=\tilde{{\mathcal{L}}}_{1}$ as
@@ D.2 Heat Diffusion in 1D
~$\tilde{{\mathcal{L}}}_{1}:=\frac{1}{(\Delta x)^{2}}\begin{bmatrix}-2&1&0&\dots&0&1\\ 1&-2&1&\dots&0&0\\ 0&1&-2&\dots&0&0\\ \vdots&&\ddots&&\vdots\\ 1&0&0&\dots&1&-2\\ \end{bmatrix}.$ (25)
@@ D.3 Heat Diffusion in 2D
The Laplacian matrix in two dimensions can be written in a block structure using the one-dimensional Laplacian ~$\tilde{{\mathcal{L}}}_{1}$ and appropriately sized identity matrices ${\bm{I}}$
~$\tilde{{\mathcal{L}}}_{2}=\frac{1}{(\Delta x)^{2}}\begin{bmatrix}(\Delta x)^{2}\tilde{{\mathcal{L}}}_{1}-2{\bm{I}}&{\bm{I}}&{\bm{0}}&{\bm{0}}&\dots&{\bm{0}}&{\bm{0}}&{\bm{0}}\\ {\bm{I}}&(\Delta x)^{2}\tilde{{\mathcal{L}}}_{1}-2{\bm{I}}&{\bm{I}}&{\bm{0}}&\dots&{\bm{0}}&{\bm{0}}&{\bm{0}}\\ {\bm{0}}&{\bm{I}}&(\Delta x)^{2}\tilde{{\mathcal{L}}}_{1}-2{\bm{I}}&{\bm{I}}&\dots&{\bm{0}}&{\bm{0}}&{\bm{0}}\\ \vdots&&&&\ddots&&&\vdots\\ {\bm{0}}&{\bm{0}}&{\bm{0}}&{\bm{0}}&\dots&{\bm{0}}&{\bm{I}}&(\Delta x)^{2}\tilde{{\mathcal{L}}}_{1}-2{\bm{I}}\\ \end{bmatrix}.$ (31)
@@ D.4 Heat Diffusion in 3D
$u(t,0,y,z)=0=u(t,1,y,z),\;u(t,x,0,z)=0=u(t,x,1,z),\;u(t,x,y,0)=0=u(t,x,y,1).$ (34)
We consider the unit-cube $\Omega=(0,1)^{3}$ with $N$ degrees of freedom per dimension, in total $N^{3}$.
The Laplacian matrix in three dimensions can again be written in a block structure using the two-dimensional Laplacian ~$\tilde{{\mathcal{L}}}_{2}$ and appropriately sized identity matrices ${\bm{I}}$
~$\tilde{{\mathcal{L}}}_{3}=\frac{1}{(\Delta x)^{2}}\begin{bmatrix}(\Delta x)^{2}\tilde{{\mathcal{L}}}_{2}-2{\bm{I}}&{\bm{I}}&{\bm{0}}&{\bm{0}}&\dots&{\bm{0}}&{\bm{0}}&{\bm{0}}\\ {\bm{I}}&(\Delta x)^{2}\tilde{{\mathcal{L}}}_{2}-2{\bm{I}}&{\bm{I}}&{\bm{0}}&\dots&{\bm{0}}&{\bm{0}}&{\bm{0}}\\ {\bm{0}}&{\bm{I}}&(\Delta x)^{2}\tilde{{\mathcal{L}}}_{2}-2{\bm{I}}&{\bm{I}}&\dots&{\bm{0}}&{\bm{0}}&{\bm{0}}\\ \vdots&&&&\ddots&&&\vdots\\ {\bm{0}}&{\bm{0}}&{\bm{0}}&{\bm{0}}&\dots&{\bm{0}}&{\bm{I}}&(\Delta x)^{2}\tilde{{\mathcal{L}}}_{2}-2{\bm{I}}\\ \end{bmatrix}.$ (35)
@@ D.5 Burgers in 1D
We will discretize the domain into $N+1$ grid points. Under periodic boundaries, only one of the boundary points can be eliminated. By convention, we choose the right-most point, leaving $N$ degrees of freedom (including the left-most point). The matrix associated with the discretized second derivative in one dimension now reads
@@ D.6 Burgers in 2D
The Laplacian matrix in two dimensions can be written in a block structure using the one-dimensional Laplacian ~$\tilde{{\mathcal{L}}}_{1}$ and appropriately sized identity matrices ${\bm{I}}$
@@ D.7 Coupled Navier-Stokes in 2D
The domain is again unit-square with bi-periodic boundary conditions. This is a system of partial differential equations for three unknowns, which are the two components of the velocity ${\bm{u}}$ and the scalar pressure field $p$. The pressure acts as a constraint to enforce incompressibility given by the second equation. The coupling of velocity and pressure can be challenging, but a simple way for a finite difference discretization is the staggered grid (Harlow & Welch, 1965). As depicted in figure 12(b), it also accounts for bi-periodicity by ignoring the top-most and right-most grid points but uses different locations to store the three unknowns. In this setting, there are equally many degrees of freedom per variable and direction. Each variable contributes $N^{2}$ entries. Hence, in total, there are $3N^{2}$.
@@
[…]
@@ Appendix F Experimental Setups
Validation $\text{Validation Error}=\frac{1}{|\mathcal{V}|}\sum_{i\in\mathcal{V}}\left(\frac{\|\bm{u}_{i}^{[t]}-\bm{u}_{i}^{[t],\text{r}}\|_{2}}{\|\bm{u}_{i}^{[t],\text{r}}\|_{2}}\right)^{2}$ (58)
@@
$\{\tilde{L}^{[e]}_{\text{val}}\}_{e}=\text{ema}(\{L^{[e]}_{\text{val}}\}_{e},\Delta e)$triangleright Smooth with Exponential Moving Average
$r=\tilde{L}^{[-1]}_{\text{val}}/\tilde{L}^{[-\delta]}_{\text{val}}$triangleright Relative change to $\delta$ epochs prior
@@
       $r_{c}=\tilde{L}^{[-1]}_{\text{val}}/c$triangleright Relative change to previous plateau
@@
             $c\leftarrow\tilde{L}^{[-1]}_{\text{val}}$triangleright Save current plateau
@@ F.1 Poisson Equation - Inverse Problem
We use $\theta_{\text{r}}=2.0$, and an initial guess for gradient descent $\theta_{\text{init}}=5.0$. 170 update steps are performed with a constant learning rate of 275. For PRDP, we set the control parameter values to $\tau_{\text{step}}=0.92$, $\tau_{\text{stop}}=0.98$, $\delta=2$ Training was started with $K_{0}=25$ linear solver iterations. At every refinement, it was incremented by $\Delta K=10$. A relatively high value for $\Delta K$ and a aggressive refinement strategy with relatively smaller value of $\tau_{\text{step}}$ are suitable for cases where $K_{\max}$ is significantly higher than the number of outer iterations. In this case, the physics converges at $K_{\epsilon}=600$ (which is also $K_{\text{max}}$ due to the double-convex problem). PRDP was successful at enabling 33% solver iteration savings
@@ F.2 Heat Diffusion - Autoregressive Neural Emulator Training
PRDP is controlled using the validation metric of equation 58 computed at $t=2$, and refinement is applied to the iterative linear solver $\mathcal{P}_{K}$ that steps from $u^{[1]}$ to $u^{[2]}$. For the 1D case, we use parameter values $\tau_{\text{step}}=0.98$, $\tau_{\text{stop}}=0.9$, $\delta=3$, exponential smoothing is performed over a time window of $8$. The 2D case uses fewer epochs, hence a less aggressive refinement using $\tau_{\text{step}}=0.9$ with a smoothing window of 6 was suitable. In the 3D case, which uses 100 epochs, used a less aggressive refinement with $\tau_{\text{step}}=0.97$, similar to the 1D case, along with $\tau_{\text{stop}}=0.9$, $\delta=3$, exponential smoothing is performed over a time window of $6$.
@@ F.4 Navier Stokes - Neural-Hybrid Corrector Learning
[…]
@@ G.1 PRDP Savings in terms of compute time
Neural Emulator Heat 2D $1.4\pm 0.08$ $1.1\pm 0.06$

### v2

@@ 2 Differentiating Iterative Linear Solvers
[…]
@@ 2.2 Scheduling inner iterations
Since the outer problem of equation 1 is solved iteratively, the gradient $\nabla_{\theta}L$ can be coarse at the beginning of the outer iterations and only requires the highest fidelity towards the end (Pedregosa, 2016). This fundamental idea was developed in the domain of hyperparameter optimization, and we demonstrate its application to physical inverse problems. Consider the Poisson equation, which is a prototypical elliptic partial differential equation found in many areas of science and engineering. Most discretization techniques lead to a linear system of equations with NN degrees of freedom (for an example discretization, refer to section D.1).
@@ 3 Progressively Refined Differentiable Physics
Figure 5 illustrates the basic concept of the PRDP control algorithm. The exact algorithm is detailed in pseudocode in Algorithm 4. Its implementation in a training pipeline that uses differentiable physics is represented by the should_refine function in listing 6.
def update(model, state_in, state_out, inner_iterations):
# neural network gives first time step, physics gives second time step
state_1 = model(state_in)
state_2 = physics(state_1, inner_iterations)
loss, grad = value_and_grad(loss_fn)(state_2, state_out)
model = update_model(model, grad)
return model, loss
[…]
val_loss_history = [evaluate(model, val_dataloader))]
for epoch in range(epochs):
for state_in, state_out in dataloader:
model, loss = update(model, state_in, state_out, inner_iterations)
val_loss_history.append(evaluate(model, val_dataloader))
if should_refine(val_loss_history):
inner_iterations += 1
Figure 6: A typical mixed-chain learning pipeline (figure 1) as used in our neural emulator learning experiments. Training begins with coarse physics, using, e.g., 1 solver iteration. The should_refine function applies PRDP, determining when to refine and progressively increasing the inner solver iterations during training.
@@ Appendix A Notation
  • $^{T}$: Transpose of a matrix or vector.
@@ C.2 Unrolled Differentiaion
$\bar{{\bm{A}}}^{T}\overset{\text{unroll}}{=}\sum_{k=0}^{K-1}\left(\bar{{\bm{u}}}_{h}^{[k]}\right)^{T}{\bm{J}}_{\Phi,2}|_{{\bm{u}}_{h}^{[k]}=\Phi^{k}({\bm{u}}_{h}^{[0]};{\bm{A}},{\bm{b}}_{h})},$ (15)
$\bar{{\bm{b}}}_{h}^{T}\overset{\text{unroll}}{=}\sum_{k=0}^{K-1}\left(\bar{{\bm{u}}}_{h}^{[k]}\right)^{T}{\bm{J}}_{\Phi,3}|_{{\bm{u}}_{h}^{[k]}=\Phi^{k}({\bm{u}}_{h}^{[0]};{\bm{A}},{\bm{b}}_{h})}.$ (16)
@@ D.1 Poisson Equation in 1D
For a finite difference discretization, the domain $\Omega=(0,1)$ is divided equidistantly into $N+2$ grid points, out of which two are related to the prescribed value on the boundary and can, hence, be ignored. As such, there are NN degrees of freedom that make up the discrete solution vector ${\bm{u}}_{h}\in\mathbb{R}^{N}$. Equally, the right-hand side is discretized at the same points and negated, yielding ${\bm{b}}_{h}$. At index $i$ interior to the domain, the second derivative is approximated with the three-point stencil
@@
in which the first matrix is due to the discretized Laplace operator in one dimension. We denote it $\tilde{{\mathcal{L}}}_{1}$ (the tilde is to distinguish it from the Laplacian matrices on periodic boundaries of the following sections). The subscript is to indicate the one-dimensional setting. It is tridiagonal and solely defines the system to be solved with ${\bm{A}}=\tilde{{\mathcal{L}}}_{1}$ as
@@ D.2 Heat Diffusion in 1D
$\tilde{{\mathcal{L}}}_{1}:=\frac{1}{(\Delta x)^{2}}\begin{bmatrix}-2&1&0&\dots&0&1\\ 1&-2&1&\dots&0&0\\ 0&1&-2&\dots&0&0\\ \vdots&&\ddots&&\vdots\\ 1&0&0&\dots&1&-2\\ \end{bmatrix}$. (25)
@@ D.3 Heat Diffusion in 2D
The Laplacian matrix in two dimensions can be written in a block structure using the one-dimensional Laplacian $\tilde{{\mathcal{L}}}_{1}$ and appropriately sized identity matrices $\bm{I}$}
$\tilde{{\mathcal{L}}}_{2}=\frac{1}{(\Delta x)^{2}}\begin{bmatrix}(\Delta x)^{2}\tilde{{\mathcal{L}}}_{1}-2{\bm{I}}&{\bm{I}}&{\bm{0}}&{\bm{0}}&\dots&{\bm{0}}&{\bm{0}}&{\bm{0}}\\ {\bm{I}}&(\Delta x)^{2}\tilde{{\mathcal{L}}}_{1}-2{\bm{I}}&{\bm{I}}&{\bm{0}}&\dots&{\bm{0}}&{\bm{0}}&{\bm{0}}\\ {\bm{0}}&{\bm{I}}&(\Delta x)^{2}\tilde{{\mathcal{L}}}_{1}-2{\bm{I}}&{\bm{I}}&\dots&{\bm{0}}&{\bm{0}}&{\bm{0}}\\ \vdots&&&&\ddots&&&\vdots\\ {\bm{0}}&{\bm{0}}&{\bm{0}}&{\bm{0}}&\dots&{\bm{0}}&{\bm{I}}&(\Delta x)^{2}\tilde{{\mathcal{L}}}_{1}-2{\bm{I}}\\ \end{bmatrix}$. (31)
@@ D.4 Heat Diffusion in 3D
u(t,0,y,z)=0=u(t,1,y,z),u(t,x,0,z)=0=u(t,x,1,z),u(t,x,y,0)=0=u(t,x,y,1).u(t,0,y,z)=0=u(t,1,y,z),\;u(t,x,0,z)=0=u(t,x,1,z),\;u(t,x,y,0)=0=u(t,x,y,1). (34)
We consider the unit-cube $\Omega=(0,1)^{3}$ with NN degrees of freedom per dimension, in total $N^{3}$.
The Laplacian matrix in three dimensions can again be written in a block structure using the two-dimensional Laplacian $\tilde{{\mathcal{L}}}_{2}$ and appropriately sized identity matrices $\bm{I}$}
$\tilde{{\mathcal{L}}}_{3}=\frac{1}{(\Delta x)^{2}}\begin{bmatrix}(\Delta x)^{2}\tilde{{\mathcal{L}}}_{2}-2{\bm{I}}&{\bm{I}}&{\bm{0}}&{\bm{0}}&\dots&{\bm{0}}&{\bm{0}}&{\bm{0}}\\ {\bm{I}}&(\Delta x)^{2}\tilde{{\mathcal{L}}}_{2}-2{\bm{I}}&{\bm{I}}&{\bm{0}}&\dots&{\bm{0}}&{\bm{0}}&{\bm{0}}\\ {\bm{0}}&{\bm{I}}&(\Delta x)^{2}\tilde{{\mathcal{L}}}_{2}-2{\bm{I}}&{\bm{I}}&\dots&{\bm{0}}&{\bm{0}}&{\bm{0}}\\ \vdots&&&&\ddots&&&\vdots\\ {\bm{0}}&{\bm{0}}&{\bm{0}}&{\bm{0}}&\dots&{\bm{0}}&{\bm{I}}&(\Delta x)^{2}\tilde{{\mathcal{L}}}_{2}-2{\bm{I}}\\ \end{bmatrix}$. (35)
@@ D.5 Burgers in 1D
We will discretize the domain into $N+1$ grid points. Under periodic boundaries, only one of the boundary points can be eliminated. By convention, we choose the right-most point, leaving NN degrees of freedom (including the left-most point). The matrix associated with the discretized second derivative in one dimension now reads
@@ D.6 Burgers in 2D
The Laplacian matrix in two dimensions can be written in a block structure using the one-dimensional Laplacian $\tilde{{\mathcal{L}}}_{1}$ and appropriately sized identity matrices $\bm{I}$}
@@ D.7 Coupled Navier-Stokes in 2D
The domain is again unit-square with bi-periodic boundary conditions. This is a system of partial differential equations for three unknowns, which are the two components of the velocity $\bm{u}$} and the scalar pressure field pp. The pressure acts as a constraint to enforce incompressibility given by the second equation. The coupling of velocity and pressure can be challenging, but a simple way for a finite difference discretization is the staggered grid (Harlow & Welch, 1965). As depicted in figure 13(b), it also accounts for bi-periodicity by ignoring the top-most and right-most grid points but uses different locations to store the three unknowns. In this setting, there are equally many degrees of freedom per variable and direction. Each variable contributes $N^{2}$ entries. Hence, in total, there are $3N^{2}$.
@@
$\mathbf{\times}$$\mathbf{\times}$$\mathbf{\times}$$\mathbf{\times}$$\mathbf{\times}$$\mathbf{\times}$$\mathbf{\times}$$\mathbf{\times}$$\mathbf{\times}$$\mathbf{\times}$$\mathbf{\times}$$\mathbf{\times}$$\mathbf{\times}$$\mathbf{\times}$$\mathbf{\times}$$\mathbf{\times}$
@@ Appendix F Experimental Setups
$\text{Validation Error}=\frac{1}{|\mathcal{V}|}\sum_{i\in\mathcal{V}}\left(\frac{\|\bm{u}_{i}^{[t]}-\bm{u}_{i}^{[t],\text{r}}\|_{2}}{\|\bm{u}_{i}^{[t],\text{r}}\|_{2}}\right)^{2}$ (58)
@@
$\{\tilde{L}^{[e]}_{\text{val}}\}_{e}=\text{ema}(\{L^{[e]}_{\text{val}}\}_{e},\Delta e)$ $\triangleright$ Smooth with Exponential Moving Average
$r=\tilde{L}^{[-1]}_{\text{val}}/\tilde{L}^{[-\delta]}_{\text{val}} \triangleright$ Relative change to $\delta$ epochs prior
@@
 $r_{c}=\tilde{L}^{[-1]}_{\text{val}}/c$ $\triangleright$ Relative change to previous plateau
@@
    $c\leftarrow\tilde{L}^{[-1]}_{\text{val}} \triangleright$ Save current plateau
@@ F.1 Poisson Equation - Inverse Problem
We use $\theta_{\text{r}}=2.0$, and an initial guess for gradient descent $\theta_{\text{init}}=5.0$. 170 update steps are performed with a constant learning rate of 275. For PRDP, we set the control parameter values to $\tau_{\text{step}}=0.92$, $\tau_{\text{stop}}=0.98$, $\delta=2$ . Training was started with $K_{0}=25$ linear solver iterations. At every refinement, it was incremented by $\Delta K=10$. A relatively high value for $\Delta K$ and a aggressive refinement strategy with relatively smaller value of $\tau_{\text{step}}$ are suitable for cases where $K_{\max}$ is significantly higher than the number of outer iterations. In this case, the physics converges at $K_{\epsilon}=600$ (which is also $K_{\text{max}}$ due to the double-convex problem). PRDP was successful at enabling 33% solver iteration savings
@@ F.2 Heat Diffusion - Autoregressive Neural Emulator Training
PRDP is controlled using the validation metric of equation 58 computed at $t=2$, and refinement is applied to the iterative linear solver $\mathcal{P}_{K}$ that steps from $u^{[1]}$ to $u^{[2]}$. For the 1D case, we use parameter values $\tau_{\text{step}}=0.98$, $\tau_{\text{stop}}=0.9$, $\delta=3$, exponential smoothing is performed over a time window of 88. The 2D case uses fewer epochs, hence a less aggressive refinement using $\tau_{\text{step}}=0.9$ with a smoothing window of 6 was suitable. In the 3D case, which uses 100 epochs, used a less aggressive refinement with $\tau_{\text{step}}=0.97$, similar to the 1D case, along with $\tau_{\text{stop}}=0.9$, $\delta=3$, exponential smoothing is performed over a time window of 66.
@@ F.4 Navier Stokes - Neural-Hybrid Corrector Learning
The loss function is defined as the sum of MSE errors of the first and second time-step solutions. We use the Adam optimizer from the Optax library, with a cosine learning rate decay schedule using an initialization of $10^{-3}$ and $\text{decay\_steps}=800$. We train in mini batches of 25 samples per iteration and for a total of 100 epochs, i.e., 800 update steps.
@@ G.1 PRDP Savings in terms of compute time
Neural Emulator Heat 2D $4\pm 0.08 1.1$$1\pm 0.06$